OpenStudy (anonymous):

WILL FAN!!!!!!!!!!!!!!! A zoo train ride costs $4 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 27, and the total money collected was $60. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers?

OpenStudy (anonymous):

A 11 children and 16 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60 B 16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60 C11 children and 16 adults Equation 1: a + c = 27 Equation 2: 4a − c = 60 D16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a − c = 60

Vocaloid (vocaloid):

let a = the number of adults let c = the number of children we know that the total number of children and adults is 27, so a + c = 27 the total amount of money collected is the adult ticket price*a + the child ticket price*c = 60 the adult ticket price is $4 the child's ticket price is $1 a + c = 27 4a + c = 60 solve for a and c

OpenStudy (hockeychick23):

B) 16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60 If we plug the numbers into each equation we get: 11+16=27 4(11)+16--> 44+16= 60